OpenStudy (anonymous):

The position of a particle on the x-axis at time t, t > 0, is s(t) = ln(t) with t measured in seconds and s(t) measured in feet. What is the average velocity of the particle for e ≤ t ≤ 2e? ln2 the quotient of 1 and the quantity of 3 times e the quotient of the natural logarithm of 2 and e the quotient of the natural logarithm of 2 and 2

OpenStudy (anonymous):

thank u!!!

OpenStudy (anonymous):

@ganeshie8 do you agree that it's b?

ganeshie8 (ganeshie8):

"average velocity" is same as "average rate of change"

OpenStudy (matlee):

Lol listen to ganeshi

OpenStudy (matlee):

im pretty sure he is more certified to say the answer

ganeshie8 (ganeshie8):

average velocity between t=2e and t=e : \[\large \dfrac{\ln(2e) -\ln(e)}{2e-e}\] simplify

OpenStudy (anonymous):

so what do i get as a result? :) btw ur an awesome helper! :D thank usomuch

ganeshie8 (ganeshie8):

saying me awesome wont get u an answer, work on simplifying the above fraction ;p

OpenStudy (anonymous):

idk how :(

ganeshie8 (ganeshie8):

simplifying denominator should be easy ?

OpenStudy (anonymous):

yes e(2-1)

ganeshie8 (ganeshie8):

\[\large \dfrac{\ln(2e) -\ln(e)}{2e-e}\] \[\large \dfrac{\ln(2e) -\ln(e)}{e}\]

ganeshie8 (ganeshie8):

use below log property for simplifying numerator : \[\large \ln (a) - \ln(b) = \ln(\dfrac{a}{b})\]